In the following situations determine the molecular structure of the compound. Show all your work and reasoning. The compound has a molecular formula C5H10O2. Based on the IR and NMR spectrum illustrated below, propose a reasonable structure for the compound
Ce зн зн 2H 0.5 PPM
Answer #1

PAGE NO - DATE Cs HiO2 Degree of unatration & (Qucta) - nH (D.u.) S = 205)+2 - 12 D.U. = 1 from IR spectrom | At ~3000 E-H st